what is the leghnth of the diagonal of a square with area 40.5 square inches?: what is the leghnth of the diagonal of a square with area 40.5 square inches?
Answer by stanbon(19670)

what is the length of the diagonal of a square with area 40.5 square inches?
--------------
Draw the picture.
Let each side of the square be "x".
Then x^2=40.5
Then, diagonal^2 = side^2 + side^2
diagonal^2 = 40.5 + 40.5 = 81
diagonal = 9 inches
=============
Cheers,
Stan H.